# Towards Zero-Shot Multimodal Machine Translation

### v1

@@ 5.2 Implementation details
We trained three different ZeroMMT with different sizes (600M, 1.3B and 3.3B) with NLLB (Costa-jussà et al., 2022) models as implemented in the Transformers library (Wolf et al., 2020) as the underlying MT system. For SigLIP (Zhai et al., 2023), we use ViT-B-16-SigLIP-384 trained on WebLI (Chen et al., 2023) from the timm library (Wightman, 2019). Following VGAMT (Futeral et al., 2023), we used bottleneck adapters (Houlsby et al., 2019) as implemented in the Adapters Python library (Poth et al., 2023) with a factor reduction of 8 and ReLU activation (Agarap, 2018) for each layer. The visual projector is a 1-layer neural network followed by ReLU activation projecting SigLIP (Zhai et al., 2023) embeddings towards the hidden dimension of NLLB. The image representation is then concatenated to the sequence of text embeddings. The cross-attention mechanism in the decoder of the sequence-to-sequence model can only attend to the positions of text embeddings. Similarly to VGAMT, we randomly mask 25% of the input tokens for VMLM.
@@
We train our models with a batch size of 32, the Adam optimizer (Kingma and Ba, 2015) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and learning rate of $10^{-4}$. We use $\lambda=0.1$ to balance the two training losses. All hyperparameters were selected based on the combination of the CoMMuTE validation set (see Section 3) and the English–French validation dataset of Multi30K, each score weighted equally. All our models are multilingual if not otherwise specified. We run each experiment three times with three different seeds and report average scores and standard error. It took 15 hours on one NVIDIA V100 for the 600M model and 20 hours on one NVIDIA A100 for the largest models.
@@
We evaluate MMT generation with BLEU (Papineni et al., 2002) and COMET (Rei et al., 2020). For BLEU, we use the Sacrebleu implementation (Post, 2018) with 13a tokenization for French, German and Czech and zh tokenization for Chinese. For COMET, we use the wmt22-comet-da (Rei et al., 2022) model from the XLM-R backbone (Conneau et al., 2020). The translations were obtained with beam search decoding of size 4 (Graves, 2012). Following (Futeral et al., 2023), we calculate disambiguation accuracy using CoMMuTE. More precisely, given an English sentence and an associated image, we compute the perplexities of each contrastive translations and the example score is 1 if the perplexity of the correct translation is lower than the perplexity of the contrastive one and 0 otherwise.
@@ 5.3 Results
We compare our approach to several others. Firstly, we compare to the text-only MT systems on which the ZeroMMT models are based, NLLB-600M distilled, NLLB-1.3B and NLLB-3.3B (Costa-jussà et al., 2022). We also compare against well-known fully supervised MMT systems: Gated Fusion (Wu et al., 2021), a tiny 3M-parameter Transformer trained from scratch on the M30K; VTLM (+ MMT) (Caglayan et al., 2021), a 44M-parameter MMT system first pretrained on the translation language modelling (TLM) objective (Lample and Conneau, 2019) with an additional image as input on translated captioning data [^10] We use the same translated data and tokenizer as in our approach. and then MMT-finetuned on M30K; and VGAMT (Futeral et al., 2023), a 630M-parameter MMT system (of which 13M trainable), which is an MT-fine-tuned mBART (Liu et al., 2020) transformed into an MMT system through the addition of lightweight adapters trained jointly on the MMT and VMLM objectives. VGAMT originally uses multiple types of visual input and is bilingual. Therefore, to have a comparable setup we retrain a VGAMT-like model with NLLB-600M distilled as the underlying MT model, with SIGLIP features only and in a multilingual setting. Finally, we compare to Multilingual OpenFlamingo (Futeral et al., 2024), a 3B multilingual multimodal language model pretrained on a large number of text-image pairs and interleaved documents which allow for zero-shot MMT, in a way that is comparable with what our model does.
We also compute an approximate upperbound on CoMMuTE for models trained with SIGLIP features by evaluating on NLLB-SIGLIP (Visheratin, 2023). Concretely, we compute the cosine similarity between the translation correctly associated with the image and the cosine similarity between the translation wrongly associated with the image for each instance of CoMMuTE. If the cosine similarity of the correct translation is higher than the wrong one, it is considered a correct prediction.
(a) English–Russian example.
(b) English–Czech example.
(c) English–Chinese example.
(d) English–Arabic example.
Figure 2: Examples from CoMMuTE and translations by our approach, ZeroMMT, and the NLLB distilled MT model.
Quantitative results
Tables 2 and 3 show the aggregated results on generation benchmarks and CoMMuTE respectively. Full results can be found in Appendix A. Compared to the text-only NLLB-600M distilled model, we observe that our approach results in only a small drop in performance on generation benchmarks (-0.52 BLEU points and -0.79 COMET points on average), where images are not required to translate the sentence correctly, despite not using the Multi30k training data or any fully supervised data. For the disambiguation task, Multilingual OpenFlamingo obtains the strongest CoMMuTE scores but it fails in generation as it was not specifically trained to translate. Our approach is significantly better than the random baseline ($>$55% for all languages for the smallest model, $>$61% on average for the largest model), showing that it is able to exploit images for disambiguation; results are close to VGAMT scores for Czech despite not having been trained on fully supervised data. Table 4 shows additional results on CoMMuTE but used as a traditional MMT generation benchmark. We obtain higher COMET scores than the text-only baseline NLLB on all languages for all model sizes except Czech for the smallest model. These results show that our approach is able to maintain good translation performance whilst still being able to exploit visual information for disambiguation. We shall see in Section 7 how the trade-off between translation quality and image exploitation can be controlled and how our approach can be made to match or outperform Multilingual OpenFlamingo both in translation quality (as measured by BLEU and COMET) and in image exploitation (as measured on CoMMuTe).
@@
Table 4: COMET scores on CoMMuTE (as a classic generation task). The best result for each model size is in bold.
@@
We analysed some translations of our ZeroMMT-600M model and compared it with the text-only distilled NLLB-600M model. We notice that for all languages, our model is able to exploit the image to slightly change the translation towards the correct meaning, as illustrated by Figures 2(c), 2(d), 2(a), 3(a), 3(b) and 2(b), where ambiguous parts of the translations change given the image fed to the model. In the example (2) of Figure 2(b), the model exploits the image to change the translation from trenéry ‘trainer’ to autobusy ‘bus’, which, although not correct, is closer to the reference translation. Again, in Figure 2(c), the translation is improved, with the word bass being translated as {CJK}UTF8gbsn鱼 ‘fish’ rather than as {CJK}UTF8gbsn低音 ‘bass (low tone)’. We also notice very few variations in the other areas of the translation in comparison with the NLLB translation, which means that our model correctly identifies the part to change.
(a) English–French example.
(b) English–German example.
Figure 3: Examples from CoMMuTE and translations by our approach, ZeroMMT, and the NLLB distilled MT model.
@@ 6 Ablation study
We conduct an ablation study on our ZeroMMT-600M model to analyse the impact of our two objectives on the results observed in Section 5.3. We first train a model without the VMLM objective, then a model without the KL penalty. We also test the replacement of the KL penalty with a standard auto-regressive MMT translation loss with the translated data as the ground truth, and finally we vary the KL penalty coefficient to observe the evolution of COMET and CoMMuTE scores.
@@
In Table 5, we show that with the KL penalty only, the model is not capable of exploiting visual information for translation. This is because there is no need for the model to use the input image and the model learns to ignore it. The aggregated CoMMuTE score of 50.3 is close to random guessing.
@@
Finally, we replace the KL penalty with the standard MMT objective (i.e. +MMT w/O KL in Table 5) as the objective to maintain translation quality. We observe that the MMT translation objective sees a drop of 0.64 BLEU points and 0.45 COMET points on average in comparison with the use of the KL penalty. It additionally results in an average drop in performance of 1.4 points on CoMMuTE.
@@
7 Controlling disambiguation level at inference time
@@
Table 6: Evolution of BLEU, COMET and CoMMuTE aggregated scores reached by our CFG-enabled ZeroMMT model over benchmarks and languages compared to the vanilla, CFG-free ZeroMMT model (i.e. with $\gamma=1.0$). Bold is best result for each model size. Underline is second best. The “MOF” line gives the corresponding scores for Multilingual OpenFlamingo.
We show that our method allows us to obtain an MMT system with a good trade-off between keeping strong translation quality and having the capacity to exploit visual context for disambiguation. However, some applications could require stronger disambiguation capabilities and be less reliant on translation fidelity or vice versa. Instead of retraining a model to control the trade-off between the two objectives, we instead propose to use classifier-free guidance (CFG) (Ho and Salimans, 2021; Sanchez et al., 2023) to control this trade-off without retraining or using fully supervised data. We define CFG in the context of MMT as follows:
@@
where $f_{\theta}$ is the text-only MT system, $f_{\theta,\beta}$ is the adapted MMT system, $x$ is the source sentence, $y$ is the generated sentence, $i$ is the visual input, $j$ is the token index and $\gamma$ is the CFG value that controls the guidance.
We run experiments to observe the evolution of BLEU and COMET scores on generation benchmarks and CoMMuTE scores when varying the $\gamma$ parameter. We show in Table 6 that for ZeroMMT-600M we can achieve a boost in CoMMuTE accuracy of up to 2.8 points for $\gamma=$ 1.5, while facing only a moderate drop of BLEU and COMET scores. Higher $\gamma$ values result in stronger disambiguation capabilities as shown by CoMMuTE but this comes at the expense of a drop in generation quality. CFG can therefore allow us to control the trade-off between disambiguation capabilities and translation fidelity depending on the application. Importantly, when $\gamma\geq 2.5$, ZeroMMT-600M matches or outperforms Multilingual OpenFlamingo on both translation quality (as measured by BLEU and COMET) and image exploitation (as measured on CoMMuTE) while being five times smaller. With larger models, we strongly outperform Multilingual OpenFlamingo on all metrics for different CFG values and we can obtain CoMMuTE scores up to 71.7 on average for $\gamma=$ 3.0 for ZeroMMT-3.3B.
@@ 8 Conclusion
We present ZeroMMT, a novel zero-shot MMT approach to bypass the need for fully supervised data when training MMT models. We observe good disambiguation capabilities such that ZeroMMT is able to effectively exploit images in order to translate English sentences correctly. At the same time, the approach manages to maintain a good translation performance, with only a very small drop in performance according to standard generation MMT benchmarks where images are not necessary to translate the sentence correctly. ZeroMMT allows us to extend MMT to new languages as it performs well on CoMMuTE for Russian and Arabic for which no fully supervised MMT training data is available. Moreover, we show that it is possible to control the disambiguation-generation trade-off using classifier-free guidance. It is therefore a step towards having MMT systems that cover a broader set of languages without having to rely on acquiring costly training data.
@@
While our approach allows us to exploit images for translation disambiguation as shown by the scores obtained on CoMMuTE, it is still behind the upperbound. Zero-shot disambiguation capabilities also come at the expense of a slight drop in translation quality as shown by BLEU and COMET scores. There are therefore areas for improvement even if our zero-shot approach is close to its fully supervised counterparts. It is nevertheless a step towards zero-shot multimodal machine translation.

### v2

@@ 5.2 Implementation details
We trained three different versions of ZeroMMT depending on the size of the underlying NLLB model [^7] As implemented in Transformers (Wolf et al., 2020). (Costa-jussà et al., 2022) (600M, 1.3B and 3.3B). For SigLIP (Zhai et al., 2023), we use ViT-B-16-SigLIP-384 trained on WebLI (Chen et al., 2023) from the timm library (Wightman, 2019). Following VGAMT (Futeral et al., 2023), we used bottleneck adapters (Houlsby et al., 2019) as implemented in the Adapters Python library (Poth et al., 2023) with a factor reduction of 8 and ReLU activation (Agarap, 2018) for each layer. The visual projector is a 1-layer neural network followed by ReLU activation projecting SigLIP (Zhai et al., 2023) embeddings towards the hidden dimension of NLLB. The image representation is then concatenated to the sequence of text embeddings. The cross-attention mechanism in the decoder of the model can only attend to the positions of text embeddings. Similarly to VGAMT, we randomly mask 25% of the input tokens for VMLM.
@@
We train our models with a batch size of 32, the Adam optimizer (Kingma and Ba, 2015) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and learning rate of $10^{-4}$. We use $\lambda=0.1$ to balance the two training losses. All hyperparameters were selected based on the combination of the CoMMuTE validation set (see Section 3) and the English–French validation dataset of M30K, each score weighted equally. All our models are multilingual if not otherwise specified. We run each experiment three times with three different seeds and report average scores and standard error. It took 15 hours on one NVIDIA V100 for the 600M model and 20 hours on one NVIDIA A100 for the largest models.
@@
We evaluate MMT generation with BLEU (Papineni et al., 2002) and COMET (Rei et al., 2020). For BLEU, we use the Sacrebleu implementation (Post, 2018) with 13a tokenization for French, German and Czech and zh tokenization for Chinese. For COMET, we use the wmt22-comet-da (Rei et al., 2022) model from the XLM-R backbone (Conneau et al., 2020). The translations were obtained with beam search decoding of size 4. Following (Futeral et al., 2023), we calculate disambiguation accuracy using CoMMuTE: given an English sentence and an associated image, we compute the perplexities of each contrastive translation, giving a score of 1 if the perplexity of the correct translation is lower than the perplexity of the contrastive one and 0 otherwise.
@@ 5.3 Results
We compare our approach to several others. Firstly, we compare to the text-only MT systems on which the ZeroMMT models are based, NLLB-600M distilled, NLLB-1.3B and NLLB-3.3B. We also compare against well-known fully supervised MMT systems: Gated Fusion (Wu et al., 2021), a tiny 3M-parameter Transformer (Vaswani et al., 2017) trained from scratch on M30K; VTLM (+ MMT) (Caglayan et al., 2021), a 44M-parameter MMT system first pretrained on the translation language modelling (TLM) objective (Lample and Conneau, 2019) with an additional image as input on translated captioning data (using the same translated data and tokenizer as ZeroMMT-600M) and then MMT-finetuned on M30K; and VGAMT (Futeral et al., 2023), a 630M-parameter MMT system (of which 13M trainable), which is an MT-fine-tuned mBART (Liu et al., 2020) transformed into an MMT system through the addition of lightweight adapters trained jointly on the MMT and VMLM objectives. VGAMT originally uses multiple types of visual input and is bilingual. Therefore, to have a comparable setup we retrain a VGAMT-like model with NLLB-600M distilled as the underlying MT model, with SIGLIP features only and in a multilingual setting. Finally, we compare to Multilingual OpenFlamingo (Futeral et al., 2024), a 3B multilingual multimodal language model pretrained on a large number of text-image pairs and interleaved documents which allows for zero-shot MMT in a way that is comparable with our model and M2KT-VPN (Hirasawa et al., 2023), a cross-lingual zero-shot MMT model based on a tiny Transformer.
We compute an approximate upperbound on CoMMuTE for models trained with SIGLIP features by evaluating on NLLB-SIGLIP (Visheratin, 2023). For each CoMMuTE instance, we compute the cosine similarity between the translation and (i) its associated image and (ii) the other image. If the cosine similarity of (i) is higher than (ii), it is considered a correct prediction.
@@
Table 4: COMET scores on CoMMuTE (as a translation set). The best result for each model size is in bold.
Quantitative results
Tables 2 and 3 show the results on CoMMuTE and the aggregated results on generation benchmarks not composed of ambiguous examples (i.e. images are not required) respectively. For full results, see Appendix A. Compared to the text-only NLLB-600M distilled model, our approach results in only a small drop in performance on generation benchmarks (-0.52 BLEU and -0.79 COMET on average), where images are not required to translate the sentence correctly, despite not using the M30K training data or any fully supervised data. For the disambiguation task, Multilingual OpenFlamingo obtains the strongest CoMMuTE scores but it fails in generation as it was not specifically trained to translate. Our approach is significantly better than the random baseline ($>$55% for all languages for the smallest model, $>$61% on average for the largest model), showing that it is able to exploit images for disambiguation; results are close to VGAMT scores (for similar model sizes) for Czech despite not having been trained on fully supervised data. Table 4 shows additional results on CoMMuTE but used as a traditional MMT generation benchmark. We obtain higher COMET scores than the text-only baseline NLLB on all languages for all model sizes except Czech for the smallest model. These results show that our approach is able to improve translation performance by exploiting images for disambiguation in cases of ambiguous examples without using any fully supervised data during training. We shall see in Section 7 how image exploitation can be controlled at inference time and how our approach can be made to outperform Multilingual OpenFlamingo in image exploitation (as measured on CoMMuTe).
(a) English–French example.
(b) English–Chinese example.
Figure 2: Translations of CoMMuTE by our approach, ZeroMMT-600M, and the NLLB distilled MT model.
@@
We analysed some translations of our ZeroMMT-600M model and compared them with those of the text-only distilled NLLB-600M model. Our model is able to exploit the image to slightly change the translation towards the correct meaning, as shown in Figure 2(a), where ambiguous parts of the translations change when the image is provided. In Figure 2(b), the translation is also improved; bass is translated as {CJK}UTF8gbsn鱼 ‘fish’ rather than {CJK}UTF8gbsn低音 ‘bass (low tone)’. We also notice few variations in the other areas of the translation with respect to the NLLB translation, which means that our model correctly identifies the part to change. More examples can be found in Appendix B.
Human evaluation
We conduct human evaluation between ZeroMMT-3.3B and NLLB-3.3B outputs to further confirm these results. We set up A/B testing where annotators were asked to assess which of the translations was better given the source translation and the accompanying image. We randomly sampled 100 examples from CoMMuTE to assess which model is better in cases of ambiguity in the source sentence and 100 examples from M30K test sets (equally represented) in cases where images do not provide additional information for translation. Figure 3 shows that ZeroMMT-3.3B is considered to be better than NLLB-3.3B by a large margin in cases of ambiguity. In cases where the image adds no additional information (i.e. the sentence to translate is non-ambiguous), Figure 4 shows that translations are considered the same in most cases (80%), and in the remaining cases NLLB-3.3B is considered to be slightly better than ZeroMMT-3.3B (11 vs. 9). The results of a similar analysis for Arabic and Chinese is given in Figures 6, 6 and 7. By manually looking at some examples, we noticed that the few unambiguous cases where NLLB-3.3B is considered superior to ZeroMMT-3.3B are due to hallucinations (since we add additional information in cases it is not necessary, this can occasionally occur).
Figure 3: French human evaluation on 100 examples from CoMMuTE. ‘Draw’ means both translations are exactly the same or considered of same quality.
Figure 4: French human evaluation on 100 examples from M30K test sets. ‘Draw’ means both translations are exactly the same or considered of the same quality.
Figure 5: Chinese human evaluation on 100 examples from CoMMuTE. ‘Draw’ means both translations are exactly the same or considered of same quality.
Figure 6: Chinese human evaluation on 100 examples from VATEX. ‘Draw’ means both translations are exactly the same or considered of same quality.
Figure 7: Arabic human evaluation on 100 examples from CoMMuTE. ‘Draw’ means both translations are exactly the same or considered of the same quality.
@@ 6 Ablation study
We conduct an ablation study on our ZeroMMT-600M model to analyse the impact of our two objectives. We first train a model without the VMLM objective, then a model without the KL penalty. We also test the replacement of the KL penalty with a standard auto-regressive MMT translation loss with the translated data as the ground truth, and finally we vary the KL penalty coefficient to observe the evolution of COMET and CoMMuTE scores. Additional ablation study on the choice of visual feature can be found in Appendix A.4.
@@
Table 5 shows that with the KL penalty only, the model cannot exploit visual information for translation. This is because there is no need to use the input image and the model learns to ignore it. The aggregated CoMMuTE score is close to random guessing.
@@
Finally, we replace the KL penalty with the standard MMT objective (i.e. +MMT w/O KL in Table 5) as the objective to maintain translation quality. We observe a drop of 0.64 BLEU points and 0.45 COMET points on average in comparison with the use of the KL penalty. It additionally results in an average drop of 1.4 points on CoMMuTE.
@@
7 Controlling the disambiguation level
@@
Table 6: Evolution of BLEU, COMET and CoMMuTE scores of our CFG-enabled ZeroMMT-3.3B model (aggregated over benchmarks and languages) compared to CFG-free ZeroMMT (i.e. $\gamma$ = 1.0). The best and second best results for each model size are shown typographically. “MOF” shows scores for Multilingual OpenFlamingo.
We show that our method allows us to obtain an MMT system with a good trade-off between strong translation quality on unambiguous examples (i.e. where images are not necessary to translate correctly) and the capacity to exploit visual context for disambiguation. However, some applications could require stronger disambiguation capabilities and be less reliant on translation fidelity on unambiguous cases or vice versa. Instead of retraining a model to control the trade-off between the two objectives, we instead propose to use classifier-free guidance (CFG) (Ho and Salimans, 2021; Sanchez et al., 2023) to control this trade-off at inference time. We define CFG in the context of MMT as follows:
@@
where $f_{\theta}$ is the text-only MT system, $f_{\theta,\beta}$ the adapted MMT system, $x$ and $y$ the source and generated sentence, $i$ the visual input, $j$ the token index and $\gamma$ the CFG value controlling guidance.
We analyse the evolution of BLEU and COMET scores on standard generation benchmarks (where text context is enough to translate correctly), and CoMMuTE scores when varying the $\gamma$ parameter. Table 6 shows that ZeroMMT-3.3B can achieve a boost in CoMMuTE accuracy of up to 4.2 points for $\gamma=$ 1.5, while facing only a moderate drop of BLEU and COMET scores on unambiguous generation benchmarks (which do not require images as additional context in theory). Higher $\gamma$ values result in stronger disambiguation capabilities, as shown by CoMMuTE, but this comes at the expense of a drop in generation quality on the unambiguous benchmarks. CFG can therefore allow us to control the trade-off between disambiguation capability and translation fidelity depending on the application. Importantly, we strongly outperform Multilingual OpenFlamingo on all metrics for different CFG values and we can obtain CoMMuTE scores up to 71.7 on average for $\gamma=$ 3.0. Results for ZeroMMT-600M and 1.3B can be found in Table 11 in Appendix A.2.
@@ 8 Conclusion
We present ZeroMMT, a novel zero-shot MMT approach bypassing the need for parallel multimodal data. ZeroMMT shows good disambiguation capabilities (it is able to effectively exploit images) while maintaining good translation results, with only a very small drop in performance according to standard generation benchmarks where images are not necessary for correct translation. ZeroMMT allows us to extend MMT to new language directions; we show that it performs well on the CoMMuTE test set for Russian and Arabic for which no parallel multimodal training data is available. Moreover, we show that it is possible to control the disambiguation-generation trade-off using classifier-free guidance. It is therefore a step towards having MMT systems that cover a broader set of languages without having to rely on acquiring costly training data.
@@
While our approach allows us to exploit images for translation disambiguation as shown by the scores obtained on CoMMuTE, it is still behind the upperbound. Zero-shot disambiguation capabilities also come at the expense of a slight drop in translation quality in cases where text context is enough to translate correctly as shown by BLEU and COMET scores. To fill this gap, a next step, which we leave to future work, would be to detect ambiguity in the source sentence and access the images only in those cases. Indeed, in most cases, images are not necessary to translate the English source sentence correctly. There are therefore areas for improvement even if our zero-shot approach is close to its fully supervised counterparts. It is nevertheless a step towards zero-shot multimodal machine translation and the expansion of MMT to new language pairs.
@@ A.1 Main results
A.2 Impact of $\gamma$ in CFG
Table 11 shows the impact of $\gamma$ on BLEU, COMET and CoMMuTE scores of our CFG-enabled ZeroMMT 600M and 1.3B models (aggregated over benchmarks and languages) compared to the vanilla, CFG-free ZeroMMT model (i.e. $\gamma$ = 1.0). See Section 7 for results with our ZeroMMT-3.3B model.
A.3 Ablation study - Full results
Tables 12, 13, 14 and 15 show the full results of the ablation study for all languages and all benchmarks.
A.4 Ablation study - Choice of visual representation
We additionally train several ZeroMMT-600M models with different types of visual encoder. As shown by Table 16, the type of visual encoder does not have an impact on global translation performances as BLEU and COMET scores do not vary a lot on standard benchmarks between models. However, we notice significant differences on CoMMuTE; the performance of CLIP, [^8] ViT-B-32 SIGLIP999vit_base_patch16_siglip_384 and SIGLIP large101010vit_so400m_patch14_siglip_384 visual encoders are about 1.5 points higher on average on CoMMuTE in comparison to VIT111111google/vit-base-patch16-224-in21k and ResNet-50.121212microsoft/resnet-50 This is probably due to the fact that VIT and ResNet-50 are trained on ImageNet, which limits their capacity to ImageNet classes while CLIP and SIGLIP-like visual encoders are trained on free-form image-text large datasets. However, all scores on CoMMuTE are well above random, therefore validating the method for different types of visual representation.